OpenStudy (anonymous):

DESPERATELY NEED HELP find real numbers a,b, and c so that the graph of the function y=ax^2+bx+c contains the points (1,2),(-2,7), and (0,3)

OpenStudy (campbell_st):

well substitute the points (1, 2) \[2 = a + b + c\] (-2, 7) \[7 = 4a -2b + c\] (0,3) this is handy 3 = c substitute the last equation into the other 2 and you will have 2 equations in 2 unknowns.. to solve simultaneously hope this helps.

OpenStudy (anonymous):

Well, one direct way would be to form three equations (in a,b and c) using given points and solve them ... Another way would take some time :-)